# Unexpected Benefits of Self-Modeling in Neural Systems

### v1

@@ Introduction
Here we tested networks trained on three different classification tasks: the MNIST, the CIFAR- 10, and the IMDB tasks LeCun et al. (1998); Krizhevsky and Hinton (2009); Maas et al. (2011). Our primary question was not whether the auxiliary task of self- modeling would improve performance on a classification task, but rather, whether it would decrease network complexity. We measured network complexity in two ways. First, we measured the width of the distribution of weights in the output layer, where narrower distributions implied reduced complexity. Second, we measured the local learning coefficient, or real log canonical threshold (RLCT), a local measure of complexity for a given function around its critical points Furman and Lau (2023). We hypothesized that adding self-modeling as an auxiliary task to networks would reduce both measures of network complexity.
@@ Methods
We also included a more theoretically-founded measure of model complexity to help compare variants: the local learning coefficient or real log canonical threshold (RLCT) Furman and Lau (2023). RLCT is a local measure of complexity for a given function around its critical points. Its initial use in statistical learning contexts was a measure of generalization meant to fill the role of parameter count in classical learning theory. We interpret the RLCT as effective model dimension and take the perspective that given two models with comparable performance, the one with a lower RLCT will be preferable. Lower RLCT indicates less overfitting and greater parameter efficiency. Unlike many other local measures of geometry, it has been shown to be robust to loss rescaling, rebutting a common critique of such approaches Dinh et al. (2017).
@@ Results
At least by the two measures used here (the width of the distribution of weights and the RLCT metric), adding self-modeling as an auxiliary task reduced network complexity. The finding was particularly clear for the more theoretically-founded measure of complexity, the RLCT Furman and Lau (2023). The result was consistent over a range of hidden layer sizes and auxiliary task weights. Self-modeling did not improve performance accuracy on the primary task. Instead, the key finding here is that, based on the complexity metrics, the networks became more efficient. The effect could be viewed as a form of regularization. The additional training incentive of predicting internal state caused the system to choose simpler functions to accomplish the main classification task.
@@ Discussion
In the present study, we tested the hypothesis of self-regularizing through self-modeling. We used a range of networks with different architectures that performed three different classification tasks (MNIST, CIFAR-10, and IMDB classifications), and we varied parameters including layer width and auxiliary task training weight. In all cases, we found that adding a self-modeling mechanism caused a reduction in network complexity. This reduction was observed in two ways. First, the distribution of weights was narrower when self-modeling was present. Encouragement of weight magnitudes that are clustered more tightly around smaller values is a goal of weight- norm regularization methods meant to improve the generalization ability of a trained model Krogh and Hertz (1992); Nowlan and Hinton (1992). The second measure of network complexity that we used, the RLCT, was also consistently smaller when self-modeling was present. RLCT is a local measure of complexity for a given function around its critical points Furman and Lau (2023). Lower RLCT indicates less capacity for overfitting and greater parameter efficiency. Not only were both measures of complexity reduced for networks that included self- modeling, but in most cases the reduction became more pronounced as greater emphasis was placed on the auxiliary task of self-modeling. The RLCT measure showed this effect particularly clearly. These results strongly support the hypothesis that self-modeling is more than simply a network learning to predict itself. Self-modeling has a fundamental, restructuring effect on networks, reducing complexity and increasing parameter efficiency.

### v2

@@ Introduction
Here we tested networks trained on three different classification tasks: the MNIST, the CIFAR- 10, and the IMDB tasks LeCun et al. (1998); Krizhevsky and Hinton (2009); Maas et al. (2011). Our primary question was not whether the auxiliary task of self- modeling would improve performance on a classification task, but rather, whether it would decrease network complexity. We measured network complexity in two ways. First, we measured the width of the distribution of weights in the output layer, where narrower distributions implied reduced complexity. Second, we measured the local learning coefficient, or real log canonical threshold (RLCT), a local measure of complexity for a given function around its critical points Watanabe (2009). We hypothesized that adding self-modeling as an auxiliary task to networks would reduce both measures of network complexity.
@@ Methods
We also included a more theoretically-founded measure of model complexity to help compare variants: the local learning coefficient or real log canonical threshold (RLCT) Watanabe (2009). RLCT is a local measure of complexity for a given function around its critical points. Its initial use in statistical learning contexts was a measure of generalization meant to fill the role of parameter count in classical learning theory. We interpret the RLCT as effective model dimension and take the perspective that given two models with comparable performance, the one with a lower RLCT will be preferable. Lower RLCT indicates less overfitting and greater parameter efficiency. The quantity is difficult to calculate in practice and we followed Lau Lau et al. (2023) to utilize stochastic gradient Langevin dynamics to estimate a localized variant of the RLCT. Unlike many other local measures of geometry, Furman Furman and Lau (2023) showed it to be robust against loss rescaling, rebutting a common critique of such approaches Dinh et al. (2017).
@@ Results
At least by the two measures used here (the width of the distribution of weights and the RLCT metric), adding self-modeling as an auxiliary task reduced network complexity. The finding was particularly clear for the more theoretically-founded measure of complexity, the RLCT Watanabe (2009). The result was consistent over a range of hidden layer sizes and auxiliary task weights. Self-modeling did not improve performance accuracy on the primary task. Instead, the key finding here is that, based on the complexity metrics, the networks became more efficient. The effect could be viewed as a form of regularization. The additional training incentive of predicting internal state caused the system to choose simpler functions to accomplish the main classification task.
@@ Discussion
In the present study, we tested the hypothesis of self-regularizing through self-modeling. We used a range of networks with different architectures that performed three different classification tasks (MNIST, CIFAR-10, and IMDB classifications), and we varied parameters including layer width and auxiliary task training weight. In all cases, we found that adding a self-modeling mechanism caused a reduction in network complexity. This reduction was observed in two ways. First, the distribution of weights was narrower when self-modeling was present. Encouragement of weight magnitudes that are clustered more tightly around smaller values is a goal of weight- norm regularization methods meant to improve the generalization ability of a trained model Krogh and Hertz (1992); Nowlan and Hinton (1992). The second measure of network complexity that we used, the RLCT, was also consistently smaller when self-modeling was present. RLCT is a local measure of complexity for a given function around its critical points Watanabe (2009). Lower RLCT indicates less capacity for overfitting and greater parameter efficiency. Not only were both measures of complexity reduced for networks that included self- modeling, but in most cases the reduction became more pronounced as greater emphasis was placed on the auxiliary task of self-modeling. The RLCT measure showed this effect particularly clearly. These results strongly support the hypothesis that self-modeling is more than simply a network learning to predict itself. Self-modeling has a fundamental, restructuring effect on networks, reducing complexity and increasing parameter efficiency.